# RAT [Uncaptioned image]: Bridging RNN Efficiency and Attention Accuracy in Language Modeling

### v1

@@ Abstract
Transformers have become the cornerstone of modern large-scale language models; however, their dependence on softmax attention poses a major computational bottleneck, particularly in long-context settings. In this work, rather than following prevalent approaches such as linear attention (or SSMs) and local attention, we introduce an intermediate design called RAT between recurrence and attention mechanisms. It partitions the input into chunks, applies a simple linear recurrence within each chunk to capture local dependencies, and then performs softmax attention across chunks to model long-range interactions. By adjusting the size of the chunk, RAT enables flexible trade-offs, combining the strengths of RNN and attention. Empirically, with a chunk size of 16, the RAT layer achieves a 7×7$\times$7 × improvement in training speed with 100K token sequences and 9×9$\times$9 × in generation at 4K sequence length, while maintaining similar or sometimes even better accuracy compared to standard attention. We demonstrate this by training 1.3B parameter models from scratch and performing large-scale evaluations, including short- and long-context benchmarks, as well as supervised fine-tuning (SFT). We further propose a hybrid architecture that interleaves RAT with local attention. By combining efficient long-range modeling with strong local interactions, this hybrid design not only improves inference speed and reduces cache memory usage compared to attention, but also consistently enhances performance, for example, achieving an average 1 point gain in commonsense reasoning tasks, up to 4 points on code tasks, and a 1 point Rouge-L increase in a summarization SFT task. Code is available at https://github.com/CLAIRE-Labo/RAT.
@@ 1 Introduction
Language modeling has long been dominated by Transformer-based architectures because of their remarkable performance across a wide range of tasks. However, their reliance on full self-attention leads to quadratic complexity with respect to sequence length, which limits scalability in long-context processing. This has motivated many efforts to design more efficient attention mechanisms. Some approaches focus on local attention [1, 2, 3], and use sparse attention or recurrence to capture long-range information. Another line of work investigates the duality between linear RNNs and attention [4, 5, 6, 7], by enforcing strict linearity constraints and introducing the notion of state expansion, which represents recurrent states as matrices rather than vectors, and shifting from per-dimension to per-head gating.
In contrast to them, we propose an alternative architecture. We argue that overusing attention in short-range contexts underutilizes its strengths and that such local processing can instead be handled more efficiently by lightweight structures such as recurrences. Based on this insight, we explore a design that lies between RNNs and attention. This design also avoids strict linearity assumptions on either side and adopts the classical recurrence form with per-dimension gating and no state expansion.
Figure 1: Comparison of three architectures: The circles represent the token representations and the arrows indicate the information flow for the 5th token. (a) Attention: Each token retains its representation and has full access to all previous tokens. (b) RNN: The information is progressively compressed along the sequence, and the current token only accesses the last hidden state. (c) RAT(L=2): Intra-chunk recurrence compresses local information, while inter-chunk attention allows direct access to previous chunks. Note that (a) and (b) can be viewed as special cases of RAT with $L=1$ and $L=T$, respectively.
Figure 2: RAT structure. The symbol $\sigma(\cdot)$ further applies a sigmoid function to the gates. Note that we do not use extra convolution or normalization in our design.
@@
We propose a RAT layer, which is a simple but effective temporal mixing mechanism. We divide long sequences into chunks: within each chunk, we apply recurrence for local modeling, and use softmax-based attention across chunks to access distant information directly (see Fig. 1). Recurrence efficiently captures short-range dependencies without suffering from the memory degradation issues common in long sequences, while attention over chunk-level representations enables long-range retrieval. Adjusting the size of the fragment $L$, RAT interpolates between attention (when $L=1$) and RNN (when $L=T$). To make training more efficient, we adopt a lightweight gated linear recurrence [8]. We further explore a hybrid design that interleaves RAT with sliding-window attention layers [1], allowing the two mechanisms to complement each other while remaining efficient.
In terms of efficiency, our implementation relies solely on PyTorch high-level operators without any custom CUDA kernels, yet already achieves a substantial speed-up. With a chunk size of $L=16$, the single layer latency to generate the token at position 4096 is up to 9×9$\times$9 × faster than that of vanilla attention. And as shown in Table 1, the maximum throughput of the full model improves by up to 10×10$\times$10 × over attention, and interleaving with local attention can also yield a 4×4$\times$4 × gain.
We pretrained models with different chunk sizes at the 1.3B scale using a 4K context length, and found their perplexity (PPL) falls between that of RNNs and attention. We then evaluated them in three settings: (1) zero-shot commonsense reasoning with short contexts, (2) long-context understanding on LongBench [9], and (3) SFT on two question answering (QA) and two summarization tasks. RAT with $L=16$ performs comparably to attention on most benchmarks and even surpasses it on some LongBench tasks. When interleaved with local attention, it achieves the strongest overall results across all variants while maintaining high efficiency. Our main contributions are as follows:
@@
    We propose the RAT layer as an intermediate design between RNN and attention for temporal mixing. We explore its detailed design and efficient implementation across three key scenarios, and observe significant speed.
@@
    We pretrain models at the 1.3B scale and evaluate them on 6 short-context, 14 long-context, and 4 SFT tasks, showing that RAT(L=16) performs competitively with attention.
@@
    We further investigate a hybrid variant that interleaves RAT with sliding-window attention, which yields the best overall performance—for example, +1 accuracy on commonsense reasoning tasks, +4 on code completion, +4 on a challenging QA SFT task, and +1 on a challenging summarization task—compared to full attention or its local variant.
@@
Model Throughput Pretrain Direct Evaluation SFT Val. CSR SQA Summ Code NQA1 QMSum token/sec PPL Avg. acc Avg. F1 Avg. Rouge-L Avg. EditSum F1 Rouge-L Attention 3052 7.61 56.7 18.2 19.5 23.9 61.3 23.4 RAT(L=16) 31170 7.67 55.6 19.6 20.2 17.4 60.8 23.3 Attention-SWA 4605 7.61 56.4 17.4 19.4 21.7 63.3 23.4 RAT(L=16)-SWA 13582 7.57 57.8 18.8 19.5 28.2 63.2 24.6
2 Method
2.1 Attention in temporal mixing
@@
$f(\cdot)$ denotes causal masking and the softmax function, and $f({\bm{q}}_{t}{\bm{K}}^{\top}_{:})$ are the attention weights used to aggregate ${\bm{V}}_{:}$. We use Python-style indexing in ${\bm{K}}^{\top}_{:},{\bm{V}}_{:}$ to highlight that each query is concerned with all keys and values. Although attention is widely considered a crucial component in state-of-the-art architectures for sequence modeling, its full-token access incurs a computational complexity of $\mathcal{O}(T^{2})$, where $T$ is the length of the sequence. This quadratic growth in FLOPs can lead to substantial inefficiency during training and inference. We view attention as a slow but lossless module in the mixing of information between tokens and aims to combine it with a fast but lossy component.
2.2 Recurrence in temporal mixing
RNNs exploit sequential dependencies in the data [10, 11]. As the model processes the sequence, it maintains a summary of past information in its hidden state, allowing each step to be computed solely from the previous state and the current input. This enables efficient inference and low FLOPs, in contrast to the full-sequence access required by attention. To further improve training efficiency, recent work has explored linear RNN variants [12, 13]. For simplicity, we adopt the design proposed in Feng et al. [8], which is similar to that in Smith et al. [14]. Using diagonal matrices and removing nonlinearities, the recurrence reduces to an EMA-like gating behavior, which can be efficiently implemented with the parallel scan algorithm [14]:
@@ 2.2 Recurrence in temporal mixing
where ${\bm{g}}_{t}$ and ${\bm{z}}_{t}\in\mathbb{R}^{D}$ denote the forget and output gates, respectively, computed via linear projections of the input followed by a sigmoid activation. Here, $D$ refers to the dimension of the model.
We find this simple design sufficient to achieve strong performance for short sequences. However, compressing the entire history to the current state, even with more complex recurrences, can lead to memory degradation over time [15]. As a result, RNN may struggle to retain information from distant time steps, which poses challenges for long-context tasks. Thus, in this paper, we place RNN as the fast yet lossy token mixing module.
2.3 RAT: a design between RNN and attention
Motivated by the complementary strengths of RNN and attention, we propose an intermediate design that combines their advantages: the speed of RNN and the global access of attention. Since RNN are effective for short sequences, while attention excels at modeling long-range dependencies, we reinterpret the input as a sequence of shorter chunks of sequences. A recurrent module is applied within each chunk to capture local dependencies, followed by cross-chunk attention to model global interactions efficiently, as shown in Fig. 1.
We divide a sequence of length $T$ into $C$ chunks of length $L$, such that $T=C\cdot L$. A token originally at position $t$ is re-indexed as $(c,l)$, where $c$ denotes the chunk index and $l$ the position within the chunk. Within each chunk, a forget gate ${\bm{g}}_{c,l}$ is used to recurrently aggregate the value ${\bm{v}}_{c,l}$ and key ${\bm{k}}_{c,l}$ vectors, yielding updated representations $\tilde{{\bm{v}}}_{c,l}$ and $\tilde{{\bm{k}}}_{c,l}$. For each query ${\bm{q}}_{c,l}$, we compute attention on the final key vectors $\tilde{{\bm{K}}}_{:,-1}$ from all preceding chunks, as well as the current chunk’s key $\tilde{{\bm{k}}}_{c,l}$. The causal masking function $f(\cdot)$ restricts attention to the chunks before it, followed by a softmax operation. Finally, an output gate is applied to produce the output:
@@ 2.3 RAT: a design between RNN and attention
The overall computation flow is illustrated in Fig. 2. Benefiting from the short length of each chunk, the RNN is both effective and computationally efficient, since its hidden state has sufficient capacity to encode short-range context without discarding much information. The attention applied across chunks then enables access to distant positions without requiring compression. As a result, the overall design preserves accuracy comparable to that of full softmax attention, while reducing computational cost, as inter-chunk attention operates over shorter contexts, with attention length inversely proportional to the chunk size.
We position this design as an intermediate architecture design between RNN and attention, aiming to combine their respective strengths, and refer to it as RAT. By adjusting the size of the chunk $L$, RAT interpolates between the RNN and the attention behaviors: when $L=T$, the model reduces to an RNN, since the entire sequence is treated as a single chunk; when $L=1$, it resembles the attention, since each token resides in its chunk. We also explored a reversed variant that applies attention within chunks and recurrence across chunks. Although this design also allows interpolation, it underperforms RAT with equal FLOPs, likely due to suboptimal allocation of computation to local attention. Further discussion is provided in Appendix C. In the following two paragraphs, we discuss the detailed design choices for RAT.
@@
We aim to match the parameter budget of RNN and attention. In addition to the projection of the output, which introduces the parameters $D^{2}$ in all three mechanisms, both attention and RNN use additional parameters $3D^{2}$. These correspond to projections for the query, key, and value in attention, and for the value, the forget gate, and the output gate in RNN. To maintain the exact total parameter count for RAT, which combines both mechanisms, we adopt lightweight designs. Specifically, we consider using low-rank projections for the gates ${\bm{g}}$ and ${\bm{z}}$, or sharing query and key projections across heads. Empirical results in Subsec. 3.3 show that sharing query ${\bm{q}}$ and key ${\bm{k}}$ typically improves performance by enabling more efficient parameter allocation. Importantly, this design does not reduce to single-head attention because the forget gate operates at a per-dimension level and introduces different elements into the gated key $\tilde{{\bm{k}}}$ used for the inter-chunk attention.
@@
We examine how to encode positional information in RAT. Although RNN captures position through its sequential structure, the attention is usually based on explicit encodings such as RoPE [16]. For RAT, we find that applying RoPE based on chunk indices rather than original token positions yields slightly better performance, and we use cross-chunk RoPE by default, as it may better align with inter-chunk attention. See Subsec. C.2 for a more detailed discussion, including exploration of the NoPE [17] encoding.
@@ 2.4 Efficiency
We discuss the efficient implementation of RAT for the three main scenarios in language modeling, and provide pseudocode for training or prefilling in 1 and for generation in 2.
Training or prefilling
In this setting, a sequence of tokens is processed in parallel, requiring the model to support efficient parallel computation. For the intra-chunk recurrent component in Eq. (missing) 3, we use PyTorch’s higher order operator associative scan to implement forward and backward passes with $\mathcal{O}(T)$ FLOPs. Compared to RNN, RAT applies the recurrent operation to both key and value representations, but chunking can reduce the scan depth from $\mathcal{O}(\log{T})$ to $\mathcal{O}(\log{L})$ and increase parallelism. In principle, if the chunk size $L$ is sufficiently small, even non-linear RNNs could be viable in terms of efficiency, which we leave for future work. For inter-chunk attention, each token also attends to its own chunk’s key and value, which vary due to causal masking. To handle this efficiently, we adopt an online softmax formulation [18]: we separately compute $f({\bm{q}}_{c,l}\tilde{{\bm{K}}}_{:,-1}^{\top})\tilde{{\bm{V}}}_{:,-1}$ and $f({\bm{q}}_{c,l}\tilde{{\bm{k}}}_{c,l}^{\top})\tilde{{\bm{v}}}_{c,l}$, then combine them by adjusting the softmax denominator. The first term is computed efficiently and in parallel using PyTorch’s flex attention, which supports custom causal masks and returns the softmax normalization term. The second term is implemented via a simple einsum.
Generation
In this setting, tokens are generated sequentially. The intra-chunk recurrent component can be implemented directly using the recurrence equation, as it only requires a single-step update and does not rely on parallel operators. For inter-chunk attention, standard implementations like flash attention [19] can be used without any modification, since no complex causal masking is required in the generation case.
In summary, the FLOPs per token of RAT is $\mathcal{O}(C\cdot D)$, compared to $\mathcal{O}(D)$ for RNN and $\mathcal{O}(T\cdot D)$ for full attention, where $C$ is the number of chunks, $D$ the model dimension and $T$ the sequence length. Notably, our current implementation achieves significant speed-ups in both training and generation without relying on custom CUDA or Triton kernels, though such kernels could further improve performance. Parallelism is particularly important for models with more than 7B parameters. We believe that RAT is compatible with both tensor and context parallelism. In particular, using a chunk size larger than 16 may reduce the communication overhead of ring-style attention in context parallelism.

### v3

# RAT [Uncaptioned image]: Bridging RNN Efficiency and Attention Accuracy via Chunk-based Sequence Modeling

@@ Abstract
Transformers have become the cornerstone of modern large-scale language models, but their reliance on softmax attention poses a computational bottleneck at both training and inference. Recurrent models offer high efficiency, but compressing the full sequence into a fixed-size and holistic representation can suffer from memory degradation in long contexts and limit fine-grained retrieval. To address this, we propose RAT, an intermediate design that bridges the efficiency of RNNs and capacity of attention. RAT partitions the input into chunks, applies recurrence within each chunk for local dependencies, and softmax-based attention across chunks for long-range interactions. This design mitigates memory degradation and enables direct access to distant tokens, while retaining computational efficiency. Empirically, with a chunk size of 16, the RAT block achieves a $7\times$ improvement in training speed for 100K sequence length and $9\times$ in generation at the 4K position, while maintaining similar performance compared to standard attention. We demonstrate this by training 1.3B parameter models from scratch and performing large-scale evaluations, including short- and long-context benchmarks, as well as supervised fine-tuning (SFT). We further propose a hybrid architecture that interleaves RAT with local attention. By combining efficient long-range modeling with strong local interactions, this hybrid design not only improves inference speed and reduces cache memory usage, but also consistently enhances performance and shows the overall best results. Code is available at https://github.com/CLAIRE-Labo/RAT.
@@ 1 Introduction
Language modeling has long been dominated by Transformer-based architectures due to their strong performance across a wide range of tasks. However, their reliance on full self-attention [attention] results in quadratic time and memory complexity with respect to sequence length, which limits scalability in long-context processing. This limitation has motivated a wave of recent efforts to revisit recurrent models or propose novel linear recurrent models such as state space models, linear attention methods [mamba, mamba2, linear_attention, gateddeltanet, xlstm, griffin, lru].
By comparing these architectures, we observe a key difference between recurrent models and self-attention. Recurrent approaches compress the full sequence history into fixed-size and holistic representations, which can lead to degraded memory when modeling long sequences and limit precise information retrieval. In contrast, self-attention retains full-token access and thus do not suffer from the two problems, but at the cost of heavy computation. This motivates us to explore an intermediate design that partially compresses the sequence while still maintaining global access.
@@
We propose a RAT layer, a simple yet effective temporal mixing method with a chunk-based design. It divides long sequences into chunks, applying recurrence within each chunk for local modeling and softmax-based attention across chunks for direct access to distant information (see 1(a)). Recurrence efficiently captures short-range dependencies while avoiding the memory degradation common in long sequences, whereas attention over chunk-level representations enables long-range retrieval. By adjusting the chunk size $L$, RAT interpolates between attention (when $L=1$) and RNN (when $L=T$).
(a) Comparison of three structures: Circles denote token representations and arrows show the information flow for the 5th token. (1) Attention: Each token keeps its own representation and has full access to all previous tokens. (2) RNN: Information is progressively compressed along the sequence, and the current token accesses only the last hidden state. (3) RAT(L=2): Intra-chunk recurrence compresses local information, while inter-chunk attention enables direct access to previous chunks. (1) and (2) can be viewed as RAT with $L=1$ and $L=T$, respectively.
(b) RAT structure. The symbol $\sigma(\cdot)$ further applies a sigmoid function to the gates.
To make RAT scalable and efficient, we further explore its positional encoding, parameter allocation, and efficient implementations that specifically address the causal masking problem in this chunk-based design. Moreover, we investigate a hybrid architecture that interleaves RAT with sliding-window attention [longformer, gau, transformer-xl, block-recurrent], which focuses computation on local windows. As RAT suggests that overusing attention in local contexts underutilizes its strengths, and such dependencies can be handled more efficiently via lightweight recurrence, the two approaches can be complementary.
We demonstrate that RAT is both efficient and performant. For instance, with a chunk size of $L=16$, the latency of a single temporal mixing block at position 4096 is up to $9\times$ lower than that of attention. As shown in Table 1, the full model achieves up to $10\times$ higher maximum throughput, and interleaving with local attention yields about $4\times$ improvement. We pretrain models at 1.3B scale and compare their results on 1) zero-shot commonsense reasoning with short contexts, (2) long-context understanding on LongBench [longbench], and (3) supervised fine-tuning on long-context tasks. RAT with $L=16$ performs on par with full attention in most benchmarks and even outperforms it on several LongBench tasks. When interleaved with local attention, it achieves the best overall results across all variants while maintaining high efficiency. Our main contributions are as follows:
@@
    We propose the RAT layer, a novel intermediate architecture that bridges the efficiency of recurrence and the capacity of attention. It compresses only local context while preserving global access, enabling direct retrieval and avoiding the memory degradation caused by full sequence compression in long contexts.
@@
    RAT is simple, scalable, and efficient. It requires no custom CUDA kernels and is naturally compatible with existing multi-dimension parallelism schemes. We also introduce a hybrid variant by interleaving RAT with sliding-window attention, enabling efficient long-range modeling with strong local interactions.
@@
    We validate RAT through extensive experiments at the 1.3B scale across diverse tasks, including 7 short-context reasoning benchmarks, 11 long-context tasks, 4 supervised fine-tuning objectives, and 9 retrieval-heavy synthetic evaluations. RAT with L=16L{=}16 shows comparable performance to full attention with $9\times$ faster single-layer decoding. Its hybrid variant with local attention yields the best overall results, along with a $3-4\times$ maximum throughput boost—for instance, +1 accuracy on commonsense reasoning, +4 on code completion, +4 on a challenging QA task, and +1 on a difficult summarization task.
@@
Model Throughput Pretrain Direct Evaluation SFT Val. CSR SQA Summ Code NQA1 QMSum token/sec PPL Avg. acc Avg. F1 Avg. Rouge-L Avg. EditSum F1 Rouge-L Attention 3052 7.61 56.9 18.2 19.5 23.9 61.3 23.4 RAT(L=16) 31170 7.67 56.7 19.6 20.2 17.4 60.8 23.3 Attention-SWA 4605 7.61 57.1 17.4 19.4 21.7 63.3 23.4 RAT(L=16)-SWA 13582 7.57 58.0 18.8 19.5 28.2 63.2 24.6
2 Overall architecture
To motivate our design, we first examine how attention and recurrence compress contextual information at the token level, along with the equations that will be reused later. We then introduce RAT as an intermediate mechanism that inherits the respective advantages of both.
2.1 Attention: Full-Token Access
@@
$f(\cdot)$ denotes causal masking and the softmax function, and $f({\bm{q}}_{t}{\bm{K}}^{\top}_{:})$ represents the attention weights used to aggregate $\bm{V}$$_{:}$. We adopt Python-style indexing in ${\bm{K}}^{\top}_{:},{\bm{V}}_{:}$ to emphasize that each query attends to all keys and values. This full-token access makes attention a dominant architecture in sequence modeling, but it suffers from high computational cost in both training and inference.
2.2 Recurrence: Full-Sequence Compression
Recurrent models maintain a summary of past information in a fixed-size representation, allowing each step to depend only on the previous state and the current input [gru, lstm]. To initiate the design of RAT, we adopt a simple and fast linear recurrence [minrnn, simple_linear_rnn, s5], which uses diagonal matrices and no nonlinearities to simplify the classic recurrence into an EMA-like gating mechanism:
@@ 2.2 Recurrence: Full-Sequence Compression
where ${\bm{g}}_{t}$ and ${\bm{z}}_{t}\in\mathbb{R}^{D}$ denote the per-dimension forget and output gates, respectively, computed via linear projections of the input followed by a sigmoid activation. Here, $D$ refers to the dimension of the model. This can be efficiently implemented using the parallel scan algorithm [s5]. We take this minimal choice to highlight the core idea, but the recurrence in RAT is not limited to this certain form and can be extended to more advanced variants such as 2D recurrence or nonlinear RNNs. We leave this as future work.
The state-based formulation of recurrence yields low computational cost and efficient inference, and performs well on short sequences. However, compressing the entire sequence history into a fixed-size and holistic state, even with more expressive designs, can still suffer from memory degradation when the sequence length grows [rnn_difficulty], and limits precise information retrieval, particularly in noisy contexts.
2.3 RAT: Chunk-Based Intermediate Design
To bridge the strong performance of attention enabled by full-token access with the efficiency of RNNs derived from full-sequence compression, we propose an intermediate design by reinterpreting the input as a sequence of shorter chunks. A recurrent module is applied within each chunk to model local dependencies, followed by cross-chunk attention to enable global interactions, as illustrated in 1(a). This design mitigates the fixed-size representation limitation of RNNs and the inefficiency of attention.
Technically, we divide a sequence of length $T$ into $C$ chunks of length $L$, such that $T=C\cdot L$. A token originally at position $t$ is re-indexed as $(c,l)$, where $c$ denotes the chunk index and $l$ the position within the chunk. Within each chunk, a forget gate ${\bm{g}}_{c,l}$ is used to recurrently aggregate the value ${\bm{v}}_{c,l}$ and key ${\bm{k}}_{c,l}$ vectors, yielding updated representations $\tilde{{\bm{v}}}_{c,l}$ and $\tilde{{\bm{k}}}_{c,l}$. For each query ${\bm{q}}_{c,l}$, we compute attention on the chunk-level key and value vectors, including $\tilde{{\bm{K}}}_{:,-1}$ for all preceding chunks, and $\tilde{{\bm{k}}}_{c,l}$ for the current chunk. The causal masking function $f(\cdot)$ restricts attention to the chunks before it, followed by a softmax operation. Finally, an output gate is applied to produce the output:
@@ 2.3 RAT: Chunk-Based Intermediate Design
The overall computation flow is illustrated in 1(b). Thanks to the short length of each chunk, the recurrent module efficiently captures local dependencies without significant information loss. Cross-chunk attention then enables global access without compressing distant information. This design preserves the accuracy of softmax-based attention while reducing computation, as inter-chunk attention operates over shorter sequences with $L$ as the FLOPs reduction ratio. We also experimented with a reversed variant by applying attention within chunks and recurrence across them, but found that standard RAT achieves better FLOPs utilization and overall performance (see Appendix C).
Benefits of the chunk design
We position this as an intermediate architecture between RNNs and attention, and refer to it as RAT. By adjusting the chunk size $L$, RAT interpolates between their behaviors: when $L=T$, it reduces to a pure RNN; when $L=1$, it resembles full attention. This allows us to pursue the trade-off between attention and RNN with a single-layer design, and offers greater flexibility for hybrid modeling by using different chunk sizes to emulate varying behaviors. From a mechanistic perspective, unlike either classic or advanced recurrence (state space or linear attention models) that rely on fixed-size and holistic representations, our chunk-based structure enables memory capacity to scale with sequence length while maintaining a fixed FLOPs reduction ratio. Its partial compression with direct access to prior chunks ensures its superior performance on retrieval-heavy tasks.
3 Scalable and Efficient Modeling with RAT
3.1 Design details
@@
We aim to control the parameters of RAT at $4D^{2}$, given $D^{2}$ for the output projection, $3D^{2}$ for the query, key, and value projections in attention, and an additional $2D^{2}$ for the two gates in RNN. We explore lightweight alternatives by using low-rank projections for the gates $\bm{g}$} and $\bm{z}$}, or by sharing query and key projections across heads. Empirical results in Subsec. 4.2 show that sharing the query $\bm{q}$} and key $\bm{k}$} slightly outperforms using low-rank gates. Notably, this design does not collapse into single-head attention, since the forget gate operates at the per-dimension level and produces distinct gated keys $\tilde{{\bm{k}}}$ for the inter-chunk attention.
@@
We examine how to encode positional information in RAT, due to the presence of cross-chunk attention. Motivated by the fact that RNN captures positional information through its sequential structure, we find that applying positional encoding at the chunk level, rather than relying on original token positions, yields slightly better fidelity. This strategy also improves length generalization, as the number of positions requiring encoding (i.e., the number of chunks) is much smaller than the full sequence length. In our main experiments, we use RoPE [rope] based on chunk indices for inter-chunk attention. In the length generalization study (see Subsec. C.2), we further explore NoPE [nope], which yields the best overall generalization performance.
Hybrid design with Local Attention
RAT is a hierarchical architecture, but not a hybrid model, as it applies the same strategy across all tokens, layers, and heads. It is compatible with various hybrid strategies and offers more flexibilities in hybrid modeling by varying the chunk size in different layers or heads. In particular, we explore interleaving RAT with sliding-window attention (SWA) [longformer, gau, transformer-xl], a widely adopted technique in recent models [cohere_llm, samba, gateddeltanet, griffin]. We find that the two are highly complementary: while local attention methods allocate most computation within fixed windows, RAT reserves attention for global access and handles local modeling more efficiently. Interleaving them enables the model to efficiently and effectively capture both short-range and long-range dependencies.
@@ 3.2 Efficiency
We discuss efficiency-related aspects of RAT, and show that our current implementation does not rely on custom CUDA or Triton kernels, yet achieves significant speed-ups in both training and generation in experiments. The pseudocode for training is provided in 1, and the decoding algorithm is shown in 2.
To begin, the FLOPs per token of RAT are $\mathcal{O}(C\cdot D)$, compared to $\mathcal{O}(D)$ for RNN and $\mathcal{O}(T\cdot D)$ for full attention, where $C$ is the number of chunks, $D$ the model dimension, and $T$ the sequence length. Most components of RAT are simple and easy to implement, except that the chunk-based design introduces a non-trivial causal masking challenge during training, which we address below.
Causal masking problem
In training, where tokens are processed in parallel, special care must be taken to apply causal masking in inter-chunk attention. First, a block-wise causal mask is required. Second, each token must also attend to its own chunk’s key and value, which should be gated up to its position. They vary across tokens due to causal masking, preventing efficient parallel computation. To address this, we adopt an online softmax formulation [onlinesoftmax]: we separately compute $f({\bm{q}}_{c,l}\tilde{{\bm{K}}}_{:,-1}^{\top})\tilde{{\bm{V}}}_{:,-1}$ and $f({\bm{q}}_{c,l}\tilde{{\bm{k}}}_{c,l}^{\top})\tilde{{\bm{v}}}_{c,l}$, and then combine the results by adjusting the softmax denominator. The first term can be implemented in parallel using existing attention frameworks, while the second is handled via a simple einsum.
Practical implementation
For training, we implement intra-chunk recurrence in Eq. () 3 using PyTorch’s associative scan, enabling forward and backward passes with $\mathcal{O}(T)$ FLOPs. Compared to full-sequence recurrence, chunking reduces scan depth from $\mathcal{O}(\log T)$ to $\mathcal{O}(\log L)$ and thus improves parallelism. For inter-chunk attention, we use PyTorch’s flex attention to implement the first term above. It supports flash attention with flexible features such as custom masks and returning the softmax denominator, and thus aligns well with our needs. For decoding, tokens are generated sequentially. Intra-chunk recurrence only requires single-step updates and can be implemented directly. For inter-chunk attention, standard implementations like flash attention [flashattention] can be used without modification, as no complex causal masking is required at inference time.
Parallelism
We think RAT is compatible with both tensor parallelism and context parallelism, which are commonly used for large model dimensions and long sequence lengths, respectively. Both the recurrence and cross-chunk attention in RAT are head-independent, making it easy to apply standard tensor parallelism by assigning different heads to different GPUs. For context parallelism, the intra-chunk recurrence is chunk-independent, allowing chunks to be distributed across GPUs. Since RAT stores much fewer number of chunk-level key/value vectors (e.g., 16× fewer than full attention), the cross-chunk attention may even avoid ring-style communication.